OpenStudy (anonymous):

You are using 0.05 gallons per mile of a SUV gas tank that holds 16 gallons. Make a T-table for the equation for this problem, using the x-values 0, 160, 320. Type in the ordered pairs you get as your answer.

OpenStudy (anonymous):

please help me my deadlines tomorrow.. if not i wont get crdeits

OpenStudy (rogue):

This is another linear equation. Let Y = # of gallons in car at miles x. m = -0.05 gallons/mile. x = # of miles. b = initial amount of gallons in the car = 16 gallons. Y = -0.05x + 16 Y (0) = 16 gallons Y (160) = -0.05*160 + 16 = 8 gallons Y (320) = -0.05*320 + 16 = 0 gallons